Jemima has three boxes whose total weight is 51 1


5 kg. Box B weighs 5 1


2 kg more than box A and box C


weighs 3 1


5 kg more than box B. Choose the correct statement.


[1] Box A weighs 17 5


6 kg.


[2] Box C weighs 21 1


30 kg.


[3] Box A weighs 42 1


2 kg.


[4] Box B weighs 12 1


3 kg


1
Expert's answer
2022-04-18T00:38:15-0400

"B=A+5\frac{1}{2}"

"C=B+3\frac{1}{5}"

"A+B+C=51\frac{1}{5}"


"B-5\frac{1}{2}+B+B+3\frac{1}{5}=51\frac{1}{5}"

"3B=53.5"


"B=17.83" kg

"A=17.83-5.5=12.33" kg

"C=17.83+3.2=21.03" kg


Answer: [2] Box C weighs "21.03" kg